the average life span of a neutron as a free particle at rest is 15 meter . it disintegrate spontaneously into an electron, proton and neutron. what is the average minimum velocity with which a neutron must leave the sun in other to reach the earth before breaking up. take the distance of earth from sun as 11*10^10?
1
Expert's answer
2018-03-08T10:20:07-0500
Total life span: t=15min=900sec,
Velocity required: v=L/t=11*〖10〗^10/900=1.22*〖10〗^8 m/s≈0.4c ,c-speed of light